OpenStudy (anonymous):

Choose the Standard Form of the equation y = −1/5x − 5.

OpenStudy (anonymous):

Nope. 5y = - x - 25 => 5y + x + 25 = 0

OpenStudy (phi):

See http://en.wikipedia.org/wiki/Linear_equation#Standard_form As you can see, there are many forms (arrangements of the equation) Standard form for this equation is \[ x+5y= -25 \] which matches Ax+By= C A , B and C are integers A is positive.